A facility has been designed to hold up to 75,000 pallets. In terms of storage, the facility currently is 90% full, on average. They ship full pallet quantities to their customers, at a constant rate of 25 pallets per hour. Assume 250 work days/year and two 8-hour shifts per day. Determine the facility’s inventory turn per year
Answer #1

Average inventory = 75000*90% = 67,500 pallets

Annual shipments = 25 pallets per hour * 8 hours per day * 250 days per year

= 50,000 pallets

Inventory turn per year = Annual shipments / Average inventory

= 50000/67500

= 0.7407
Answer #2

To determine the facility's inventory turnover per year, we need to calculate the average inventory and divide it by the annual shipments. Here are the steps to calculate it:

Step 1: Calculate the average inventory:

  • The facility is currently 90% full on average, which means it has an average inventory of 0.9 * 75,000 = 67,500 pallets.

Step 2: Calculate the annual shipments:

  • The facility ships at a constant rate of 25 pallets per hour.

  • There are 2 shifts per day, each with 8 hours of work.

  • So, the facility operates for 2 * 8 = 16 hours per day.

  • The facility operates for 250 work days per year.

  • The annual shipments can be calculated as follows: Annual Shipments = (25 pallets/hour) * (16 hours/day) * (250 days/year)

Step 3: Calculate the inventory turnover:

  • The inventory turnover is the ratio of annual shipments to average inventory.

  • Inventory Turnover = Annual Shipments / Average Inventory

Let's calculate the inventory turnover using the given values:

Average Inventory = 67,500 pallets Annual Shipments = (25 pallets/hour) * (16 hours/day) * (250 days/year)

Inventory Turnover = Annual Shipments / Average Inventory

Substituting the values: Inventory Turnover = [(25 pallets/hour) * (16 hours/day) * (250 days/year)] / 67,500 pallets

Simplify and calculate the inventory turnover.


answered by: mervetokaz